planet 1 is traveling at .5c relative to planet 2.   on planet 1 there is a 10 cm diameter clock. How wide is the clock and how long does it take the minute hand to go around once, in accordance to planet 2?
Answer #1

V = 0.5c and d = 10cm , t = 1mint = 60sec.

From relativistic formula

d' = d (1 - v^2/c^2)^-1/2 = (10cm) ( 1 - 0.25)^1/2 = 8.66cm .

That means clock diameter in accordance of planet 2 is d' = 8.66cm.

Similarly , t' = t/(1 - v^2/c^2)^1/2 = 60sec / 0.866 = 69.2sec.